You manage a bond portfolio with a current value of $150,000,000 & a duration of 7.32. You need to hedge the interest rate risk of this portfolio for some reason. Today's date is Monday 12/10/2018, so the settlement price for a treasury bond is the 11th. You decide to use the 10 year t-note futures to hedge. The cheapest to deliver bond is the 3 percent coupon bond with maturity date of 09/30/2025 which is currently selling for a yield of 2.766% (price of 101.4766 or 1.014766 per $1 of face value).

Question: Do you long or short these contracts?
Answer #1
STEP 1)Cash market position=long S&P 500 shares worth=$ 6000000
STEP 2)futures market value=beta*cash market value=2.4*6000000=$ 14400000
we have to sell futures worth $ 14400000
STEP 3) Number of contracts=
number of nifties=14400000/2250=6400
since 1 contract has 100 futres so 64 contracts should short i.e sell
STEP 4)hedging strategy is sell 64 contracts at 2250
fair futures price=spot price*e^t(r-d)
250*e^1(1%-2%)=250*e^-0.01
e^0.01=1+0.01/1+(0.01)^2/2+(0.01)^3/3=1.01005
e^-0.01=1/1.01005=7.43
fair furures price=250*1.01005=252.51
on 13 month
cash market futures market
buy:250 sell:2250
sell:252.51 buy:2300
profit=2.51 loss:50
total profit=2.51*24000=60240 total loss=50*64 contracts*100=320000
(6000000/250=24000)
over all loss=320000-60240=259760 $
if drops to 1200
cash market futures market
buy:250 sell:1200
sell:252.51 buy:1205
profit=2.51 loss:5
total profit=2.51*24000=60240 total loss=5*6400=32000
(6000000/250=24000)
overall profit=60240-32000=28240
if drops to 1207
cash market futures market
buy:250 sell:2100
sell:252.51 buy:2107
profit=2.51 loss:7
total profit=2.51*24000=60240 total loss=7*6400=44800
(6000000/250=24000)
overall profit=60240-44800=15440
if increases to 3200
cash market futures market
buy:250 sell:3200
sell:252.51 buy:3215
profit=2.51 loss:15
total profit=2.51*24000=60240 total loss=15*6400=96000
(6000000/250=24000)
overall loss=96000-60240=35760